HUBBLE SPACE TELESCOPE DAILY REPORT #971 PERIOD COVERED: 0800 EST Daily Status Report as of 257/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.2 Completed FOS/BL 4939 (Quasar Absorption Line Survey: Cycle 2 Observations) The Faint Object Spectrograph was used to make ultraviolet spectral observations of quasar 3C249-1 (which is also an active radio source). The observations were made in the rapid mode using grating G130H. This object is replacing quasar IR0450-2958 in proposal 3791, which could not be acquired with a binary search. A total of six observations were made. The first three observations were reported that the target may not have been well centered. The final three observations were reported to look fine. 1.3 Completed WFC 4381 (HST Medium - Deep Survey - Cycle 3) The Wide Field Camera was used to make three sets of medium-deep high latitude sky survey observations taken in parallel with the prime science observations. The observations were executed as scheduled, and no problems were reported. The image were reported to look fine. 1.4 Completed PC 4140 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor - Part 2) The WFPC was used to make internal flat calibrations images to check for new forms of contamination (i.e., Daises and measles), and to check for the possible return of quantum efficiency hysteresis. The observations completed as scheduled, and no problems were reported. 1.5 Completed FOC/96 3747 (High Resolution Observations of Cataclysmic Variable - Cycle 2) The Faint Object Camera was used to make observations of the cataclysmic variable star RR-Telescopium. This star system is suspected to posses a complex of surrounding emission nebulosity which was studied using narrow band filters centered on ultraviolet emission features of magnesium, carbon, and oxygen. The observations were executed as schedule, and no problems were reported. 1.6 Completed GHRS/2 3472 (The Properties of Single Interstellar Clouds - Cycle 2) The High Resolution Spectrograph was used to make Echelle spectral observations of the broad-lined, early-type star HD143018. Both the Echelle and grating G160M were used. All of the observations were executed, received and displayed without problems. 1.7 Completed WFC/PC 5026 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor) The Wide Field Camera and the Planetary Camera were used to make internal flat calibration images. These images will be used to look for measles and daisies (i.e., different forms of contamination), and to monitor scattered light. The observations were executed as scheduled, and no problems were reported. 1.8 Completed S/C/HSP/PMT 5079 (Dynamics of Planetary Upper Atmospheres - Cycle 3 High) The High Speed Photometer was used to make observations of star U111 which occulted the planet last April. The observations lasted about 40 minutes. FGS telemetry was also obtained via FN format, in order to monitor the pointing stability. The OSS monitored the observations and assessed the data quality. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 7 Successful : 7 2.3 Significant HSTARs: None 2.4 Operations Note: The Internal Simulation #5 - Repeat 2 (Deploy) is currently under way. The simulation should last for about 12 hours. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete PC 5073 (WF/PC Cycle 3 Calibration: PC Flats Wide. 2) 3.2 Complete Two Sets of FGS 5100 (HIPPARCOS Plate Scale Calibration: Cycle 3 Instrument Scientist Calibration) 3.3 Complete PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) 3.4 Complete GHRS/2 4873 (Local Interstellar Medium and D/H Ratio - Cycle 3 Baseline Time - Medium) 3.5 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.6 Complete Three Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.7 Complete WFC 5075 (WF/PC Cycle 2 Calibration: WFC Flats Wide. 1) 3.8 Complete FOS 4549 (Ultraviolet Spectropolarimetry of Comet P/West-Kohoutek-Ikemmmura) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-